# Towards Inducing Document-Level Abilities in Standard Multilingual Neural Machine Translation Models

### v2

@@ 5.2 Qualitative Analysis
In the preceding section, we conducted a string-metrics-based evaluation on standard benchmarks such as IN22-Conv, ALT, and FLORES, covering both sentence-level and document-level variants. However, the effectiveness of these string-based metrics in evaluating translation quality, especially in longer-context scenarios where factors like semantic coherence, coreference resolution, and continuity are critical, remains underexplored.
@@ 6 Conclusion
In this study, we have demonstrated that sinusoidal Positional Embeddings can be efficiently replaced with ROPE or ALIBI in a standard transformer model. We show that this substitution can be achieved in a parameter-efficient manner by simply introducing additional parameters within the self-attention module. Although our experiments are focused on the machine translation (MT) task, we believe that this approach can be extended to other Encoder-Decoder models or Encoder-only models to enhance their long-context capabilities.
@@ 8 Acknowledgements
We thank Jay Gala (Research Assistant, MBZUAI), and Raj Dabre (Senior Researcher, NICT) for their invaluable insights and comments on paper. We also thank Amit Sharma (Principal Researcher, Microsoft Research India) for the discussions that helped us in refining the idea.
@@ A.1 Scores
The raw scores corresponding to fig. 2 and fig. 3 shown in Table table 5 and table 6.
@@
DATASET IT2 SINE ALIBI ROPE SINE ALIBI ROPE SINE ALIBI ROPE
IN22 Gen 48.4 47.7 47.6 47.8 47.0 47.1 46.9 47.1 46.9 47.1
IN22 Conv 44.6 43.6 43.8 43.8 42.9 43.1 42.9 43.0 43.1 43.1
En-Indic FLORES 47.8 48.3 48.3 48.2 47.8 47.7 47.7 47.7 47.6 47.6
IN22 Gen 62.0 62.0 61.2 61.8 61.6 61.0 61.4 61.6 60.9 61.5
IN22 Conv 55.7 55.3 55.0 55.1 55.4 54.7 55.3 55.5 55.2 55.4
Indic-En FLORES 59.5 59.7 59.2 59.4 59.4 58.8 59.2 59.3 58.8 59.1
Table 5: ChrF++ scores on sentence-level evaluation benchmarks.
FT LoRA min LoRA
DATASET IT2 SINE ALIBI ROPE SINE ALIBI ROPE SINE ALIBI ROPE
ALT 55.0 55.2 55.4 55.3 55.1 55.0 55.0 55.0 55.0 54.9
IN22 Conv 32.4 36.2 42.4 45.1 29.6 33.5 38.6 29.3 34.1 39.4
En-Indic FLORES 49.7 50.3 49.9 50.2 49.6 48.9 49.3 49.3 48.5 49.2
ALT 65.2 65.2 64.8 65.1 65.2 64.9 65.0 65.2 64.9 64.9
IN22 Conv 32.4 40.3 51.8 56.6 38.8 47.1 55.1 35.3 43.8 53.7
Indic-En FLORES 61.2 61.2 60.0 61.2 61.1 59.4 60.9 61.2 59.1 60.9
Table 6: ChrF++ scores on document-level evaluation benchmarks.
@@ A.2 Qualitative Examples
The MQM annotations generated by GPT-4 for a test conversation across for each baseline is presented in Tables 7, 8 and 9.

### v3

@@ 5.2 Qualitative Analysis
In the preceding section, we conducted a string-metrics-based evaluation on standard benchmarks such as IN22-Conv, ALT, and Flores, covering both sentence-level and document-level variants. However, the effectiveness of these string-based metrics in evaluating translation quality, especially in longer-context scenarios where factors like semantic coherence, co-reference resolution, and continuity are critical, remains under-explored.
@@
6 Additional Experiments
To assess the scalability of our approach, we extend our methodology to the larger 1B IndicTrans2 model. For consistency, we reuse the same training data and hyperparameters as described in Section 3.2.1 and Table 4, respectively. In this case, we focus solely on FFT, as it yielded the highest performance recovery among the fine-tuning strategies explored in Section 3.4. Table 5 presents the results of this experiment. Consistent with our previous findings, RoPE continues to exhibit the best length generalization, followed by ALiBi, as observed on the Document-Level IN22-Conv subsset.
FT
DATASET IT2 Sine ALiBi RoPE
IN22-Gen 63.8 63.6 62.8 63.3
IN22-Conv 56.1 56.7 55.0 55.3
Indic-En Flores 61.1 61.0 60.2 60.5
IN22-Gen 48.3 48.2 48.4 48.7
IN22-Conv 44.8 45.3 45.9 45.9
En-Indic Flores 47.8 48.3 48.3 48.3
ALT 66.9 66.8 66.5 66.7
IN22-Conv 31.7 40.8 51.0 58.3
Indic-En Flores 62.8 62.6 61.7 62.5
ALT 55.5 55.9 55.8 55.9
IN22-Conv 28.6 35.5 45.3 48.6
En-Indic Flores 49.1 49.8 49.5 49.8
Table 5: IT2-1B ChrF++ scores on sentence-level (top) and document-level (bottom) evaluation benchmarks. The best score is bolded and second best is underlined.
@@ 7 Conclusion
In this study, we demonstrated that standard NMT systems can be efficiently retrofitted with relative positional embeddings such as RoPE and ALiBi, enabling them to handle longer contexts. Moreover, we showed that this retrofitting can be achieved through parameter-efficient fine-tuning using minimal document-level data, even across a few languages, with evidence of cross-lingual transfer. Although our experiments focused on the machine translation (MT) task, we believe that this approach can generalize to other encoder-decoder and encoder-only models, potentially improving their long-context capabilities in a wide range of applications.
@@ 9 Acknowledgements
We thank Raj Dabre (Senior Researcher, NICT), Jay Gala (Research Assistant, MBZUAI) and Amit Sharma (Principal Researcher, MSRI) for their for the discussions that helped us in refining the idea.
@@ A.1 Scores
The raw scores corresponding to Figure 2 are shown in Table 6.
@@
DATASET IT2-Dist Sine ALiBi RoPE Sine ALiBi RoPE Sine ALiBi RoPE
IN22-Gen 48.0 48.2 48.7 48.8 47.9 47.8 47.7 47.7 47.3 47.6
IN22-Conv 44.6 45.5 45.7 45.5 44.1 44.4 44.3 44.0 44.1 44.1
En-Indic Flores 47.7 48.4 48.5 48.5 47.9 47.8 47.9 47.8 47.7 47.7
IN22-Gen 62.0 62.2 61.8 62.1 61.8 61.1 61.5 61.7 61.1 61.5
IN22-Conv 55.7 55.8 55.5 55.6 55.7 55.1 55.3 55.8 55.4 55.6
Indic-En Flores 59.5 59.7 59.4 59.6 59.4 58.9 59.1 59.4 59.0 59.2
ALT 55.0 55.2 55.2 55.4 55.1 55.0 54.8 55.2 55.0 55.0
IN22-Conv 33.7 38.9 44.1 47.3 31.6 37.1 39.3 30.9 36.2 40.4
En-Indic Flores 49.0 49.7 49.9 50.4 49.7 49.0 49.5 49.6 48.9 49.5
ALT 65.2 65.3 64.8 65.1 65.2 65.0 65.0 65.2 64.9 65.0
IN22-Conv 33.8 40.1 52.6 57.1 38.4 47.3 55.0 32.0 46.3 54.6
Indic-En Flores 61.2 61.4 60.0 61.0 61.2 59.6 60.9 61.3 59.5 61.1
Table 6: IT2-Dist-200M ChrF++ scores on sentence-level (top) and document-level (bottom) evaluation benchmarks. The best score is bolded and second best is underlined.
@@ A.2 Qualitative Examples
The MQM annotations generated by GPT-4o for a test conversation across for each baseline is presented in Tables 7, 8 and 9.